A rectangular goat pen has a perimeter of 24 meters and an area of 32 square meters. What are the dimensions of the pen?

____meters by _____meters

Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52787)

.
When you leave blank spaces for me to fill them with the answer,


I understand it in the way that you are interested in getting the answer only and are not interested in learning the solution.


    OK.


The answer is  8 meters by 4 meters.

-----------------

Please read my advise very attentively.

If you are interested  TO  LEARN  from this site,  NEVER  place these blank spaces:  the tutors  HATE  it.


It is one of the basic (and self-evident) rules for visitors at this forum.



Answer by stanbon(75887)

A rectangular goat pen has a perimeter of 24 meters and an area of 32 square meters. What are the dimensions of the pen?
Equations:
2(L+W) = 24 meters
L*W = 32 meters^2
===================
Modify::
L + W = 12 m
----
Substitute to get::
L(12-L) = 32
-L^2+12L-32 = 0
L^2 - 12L +32 = 0
(L-8)(L-4) = 0
----

_8___meters by _4____meters
Cheers,
Stan H.